OpenStudy (anonymous):

Solve. -21.536=-1.6(9-m) Enter only the value of the variable.

mathslover (mathslover):

\[-21.536/-1.6 = 9-m\] \[21.536/1.6 = 9-m\] \[m = 9-21.536/1.6\] \[m=-4.46\]

OpenStudy (anonymous):